# QMetro++ - Python optimization package for large scale quantum metrology with customized strategy structures

### v1

@@ 3.2 Single channel QFI optimization
7iss_qfi, qfis, rho0, sld, status = iss_channel_qfi(channel, aniclla_dim)
@@
Fig. 4(B) and Fig. 4(C) present plots of the results obtained using iss_channel_qfi and mop_channel_qfi for the two decoherence models respectively. Since MOP method gives the optimal QFI implicitly assuming arbitrary ancilla dimension $d_{\mathcal{A}}$, these plots indicate that for the dephasing case (B) ancilla is in fact unnecessary, while in the amplitude damping case (C) ancillary system of dimension $d_{\mathcal{A}}=2$ is required for an optimal precision.
This example clearly shows the benefits of having two independent optimization procedures. One method (MOP) provides the true optimal QFI, but does not allow us to get an insight into the required ancilla size, while the other one (ISS) allows us to study the impact of the ancilla size.
@@ 3.3 Parallel strategy optimization
Figure 5: (A) Diagram of parallel strategy with multiple parametrized channels $\Lambda_{\theta}$ and ancilla $\mathcal{A}$. Values of QFI normalized by number of channel uses - $N$ for: (B) dephasing ($p=0.75$) (C) amplitude damping ($p=0.75$) and different methods: MOP - black $\times$, simple ISS with $d_{\mathcal{A}}=2$ black +, tensor network ISS with $d_{\mathcal{A}}=2,\;r_{\mathrm{MPS}}=\sqrt{r_{\mathfrak{L}}}=2$ blue dashed line, tensor network ISS with $d_{\mathcal{A}}=2,\;r_{\mathrm{MPS}}=\sqrt{r_{\mathfrak{L}}}=4$ red dotted line, upper bound - black solid line.
Parallel strategy is a strategy where $N$ copies of $\Lambda_{\theta}$ are probed in parallel by an entangled state and their output is collectively measured, see Fig. 5(A). It can be also understood as a strategy with a single channel $\Lambda_{\theta}^{(N)}=\Lambda_{\theta}^{\otimes N}$. QMetro++ package provides several functions computing QFI for parallel strategy. The simplest two, iss_parallel_qfi and mop_parallel_qfi are straightforward generalizations of ISS and MOP methods for parallel strategy:
@@
8iss_qfi, qfis, rho0, sld = iss_parallel_qfi(channel, N, aniclla_dim)
@@
This problem can be circumvented using tensor networks MPS (31) and MPO (32).This approach is implemented in iss_tnet_parallel_qfi, which requires two additional arguments specifying the size of MPS and pre-SLD MPO bond dimensions:
@@
11qfi, qfis, psis, Ls, status = iss_tnet_parallel_qfi(channel, N, aniclla_dim, mps_bond_dim, L_bond_dim)
In contrast to the function implementing the basic ISS, which returns the optimal input state and the SLD in terms of arrays, this procedure gives them in a form of lists containing tensor components: psis = [$\psi_{0}$, $\psi_{1}$, …] and Ls = [$\mathfrak{L}_{0}$, $\mathfrak{L}_{1}$, …], see Eqs. (31) and (32).
Naturally, increasing ancilla and bond dimensions allows to achieve higher values of QFI up to some optimal value, see Fig. 5(B) and (C). One might asses this value by running iss_tnet_parallel_qfi for increasing dimension sizes, but it can be also upper-bounded using the par_bounds function which gives a bound on the QFI of parallel strategy for all ancilla and bond dimension sizes (see section 2.5):
@@
The numerical results are presented in Fig. 5(B) and (C). We considered parallel strategy with ancilla dimension $d_{\mathcal{A}}$ and various bond dimensions $r_{\mathrm{MPS}}=\sqrt{r_{\mathfrak{L}}}$. Notice that for dephasing case (B) results from MOP and ISS methods coincide perfectly and that QFI for MPS with $r_{\mathrm{MPS}}=4$ approaches the upper bound with growing $N$. This suggests that in this case $d_{\mathcal{A}}=2$ is sufficient to saturate the bound. On the other hand, for the amplitude damping case (C) the results obtained using MOP and ISS differ, and already for $N=5$ the ancilla dimension $d_{\mathcal{A}}=2$ is suboptimal. Additionally, $F_{Q}^{(N)}(\Lambda_{\theta})/N$ computed using MPS quickly flattens out and the increase of bond dimension leads to a very small improvement. Therefore, we can conclude that in this case the $d_{\mathcal{A}}$ considered are insufficient to saturate the bound.
@@ 3.4 Adaptive strategy optimization
Figure 6: (A) Diagram of adaptive strategy with multiple parametrized channels $\Lambda_{\theta}$ and ancilla $\mathcal{A}$. When the optimization is performed over all quantum combs $C$ (area bounded by dashed lines) one can control the dimension of only the final ancilla ($\mathcal{A}_{\mathrm{end}}$). Values of QFI normalized by number of channel uses - $N$ for: (B) dephasing ($p=0.75$) (C) amplitude damping ($p=0.75$) and different methods: MOP - black $\times$, simple ISS with $d_{\mathcal{A}}=2$ black +, tensor network ISS with $d_{\mathcal{A}}=2$ blue dashed line, tensor network ISS with $d_{\mathcal{A}}=2$ red dotted line, upper bound - black solid line, asymptotic bound - black dash-dotted line.
In the adaptive strategy we start with an initial state of the system and the ancilla, $\rho_{0}$, and we act on the system with $N$ parameter encoding channels $\Lambda_{\theta}$, intertwined with control operations $\mathrm{C}_{i}$ which act both on the system and the ancilla, see Fig. 6(A). Equivalently, it is a strategy where $N$ parameter encoding channels $\Lambda_{\theta}$ are plugged in between the teeth of a quantum comb $C$, see Fig. 3, and then the resulting state is measured. In this strategy, one tries to obtain the value of QFI for the optimal input state, measurement and control operations (or equivalently the optimal comb and measurement). Notice that for large enough ancilla dimension $d_{\mathcal{A}}$ and for $\mathrm{C}_{i}$ being appropriate SWAP gates one can simulate any parallel strategy with an adaptive strategy. Thus for $d_{\mathcal{A}}$ large enough QFI for adaptive strategy always upper-bounds QFI for parallel strategy.
@@
Alternatively, we may also use the ISS method, with the help of iss_adaptive_qfi function. This procedure will perform the optimization over the whole comb as well. Note that, constraints on a quantum comb (19) allow for its decomposition into control operations (CPTP maps), but do not specify the ancilla dimension inside the comb. Therefore, in this case one can control only the dimension of the last ancilla which goes outside of the comb, indicated by the symbol $\mathcal{A}_{\mathrm{end}}$ in Fig. 6(A):
@@
10qfi, qfis, comb, sld, status = iss_adaptive_qfi(channel, N, aniclla_dim)
@@
Finally, if we want to make use of the tensor-network structure, and perform the optimization over separate control operations, we can use iss_tnet_adaptive_qfi. Notice that in this approach we control ancilla dimension at each step and thus, in general, the obtained QFI will be smaller then the one we got with iss_adaptive_qfi—the advantage here is that the optimization will be much more efficient (not to mention that this is also the only way to go to the larger $N$ limit):
@@
10qfi, qfis, teeth, sld, status = iss_adaptive_qfi(channel, N, aniclla_dim)
@@
The returned argument teeth represents the comb’s teeth and it is a list consisting of the optimal initial state and CJ matrices of the optimal control operations teeth $=[\rho_{0},C_{0},C_{1},...,C_{N-2}]$.
Upper bound for the adaptive strategy is computed using ad_bounds function:
@@
Additionally, one can compute the value of the bound in the limit $N\rightarrow\infty$ which is asymptotically saturable and the same for both parallel and adaptive strategy, see Sec. 2.5:
@@
6c, k = asym_scailing_qfi(channel)
@@
Values of QFI, bounds and asymptotic QFI for adaptive strategy are presented in Figs. 6 (B) and (C). In contrast to the parallel strategy, this time the amplitude damping case (C) shows results which are much closer to the bound.
Notice that there is no data for amplitude damping and ancilla dim $d_{\mathcal{A}}=4$ and $N\geq 60$. This is due to the fact that for this noise model, the QFI attains a large value and simultaneously optimal control operations $\mathrm{C}_{i}$ take the state close to the pure state. These two processes combined create objects (tensors, matrices etc.) which hold values from a very wide range at the same time what causes loss of precision and then numerical instability.
@@ 3.5 Correlated noise models
Figure 7: (A) Diagram of $N$ parametrized channels $\Lambda_{\theta}$ with noise correlations. Extended channels $\tilde{\Lambda}_{\theta}$ (gray) act on the system as $\Lambda_{\theta}$ and they are connected with each other by the environment space $\mathcal{E}$. The environment system $\mathcal{E}$ starts in the state $\rho_{E}$ and it is not accessed (it is traced out) at the end. Values of QFI normalized by number of channel uses - $N$ for anti-correlated ($c=-0.75$) dephasing ($p=0.75$), different strategies (B) parallel (C) adaptive and various methods: MOP - black $\times$, simple ISS with $d_{\mathcal{A}}=2$ black +, tensor network ISS with $d_{\mathcal{A}}=2$ (parallel: $r_{\mathrm{MPS}}=\sqrt{r_{\mathfrak{L}}}=2$) - blue dashed line, upper bound with block size $m=1$ black dotted line, upper bound with block size $m=3$ black solid line and upper bound for uncorrelated channels $c=0$ black dash-dotted line. Notice that since there are no separate bounds for parallel strategy with correlated channels the upper bounds (black dotted and solid lines) are identical in (B) and (C).
QMetro++ gives also a simple way to compute QFI for adaptive and parallel strategy in scenarios when parameter encoding channels act on an additional environment space $\mathcal{E}$ which directly connects subsequent channels, see Fig. 7(A). This approach allows in particular to model noise correlations affecting the sensing channels.
To illustrate this, let us consider a Kraus representation for the dephasing model with Kraus operators $K_{\pm}$ acting as Bloch sphere rotations by ±$\pm\epsilon$ around $z$ axis, see Eq. (44). We want the signs of these rotations to form a binary Markovian chain. The conditional probability of a rotation with sign $s_{i}\in\{+,-\}$ in channel $i$ assuming $s_{i-1}$ in channel $i-1$ is:
@@
where $c\in[-1,1]$ is a correlation parameter with $c=0$ meaning no correlations, $c=1$ maximal positive correlations and $c=-1$ maximal negative correlations. The initial probabilities are $p(\pm)=1/2$.
@@
This process can be modeled with the help of a two dimensional environment Hilbert space $\mathcal{E}$ which will carry the information about the sign of the rotation in the preceding step. Let us introduce a new channel acting on both the environment and the system $\tilde{\Lambda}_{\theta}=\Phi_{\theta}*\mathrm{T}$, see Fig. 8, where:
$\Phi_{\theta}:\mathcal{L}(\mathcal{E}\otimes\mathcal{S})\ni\rho\mapsto V_{\theta}\rho V_{\theta}^{\dagger}\in\mathcal{L}(\mathcal{E}\otimes\mathcal{S}),\\ V_{\theta}=\sqrt{2}\sum_{s=\pm}\ket{s}\bra{s}\otimes K_{s,\theta},$ (51)
@@
2from qmetro import *
3
4p = 0.75
5c = -0.75
6
7# rot_like=True gives dephasing in
8# the second form:
9Lambda = par_dephasing(p, rot_like=True)
10krauses, dkrauses = Lambda.dkrauses()
11# Rotations by plus and minus epsilon:
12Kp, Km = krauses
13# Derivatives of Krauses by theta:
14dKp, dKm = dkrauses
15
16# Plus state:
17plus = array([1, 1]) / sqrt(2)
18# Minus state:
19minus = array([1, -1]) / sqrt(2)
20
21# V_theta matrix:
22V = kron(ket_bra(plus, plus), Kp)
23V = V + kron(ket_bra(minus, minus), Km)
24V = sqrt(2) * V
25
26# Derivative of V_theta matrix:
27dV = kron(ket_bra(plus, plus), dKp)
28dV += kron(ket_bra(minus, minus), dKm)
29dV = sqrt(2) * dV
30
31# env_dim=d tells ParamChannel
32# constructor that the channel acts
33# on the d-dimensional environment space
34# and the system
35Phi = ParamChannel(

### v2

@@ 3.2 Single-channel QFI optimization
7iss_qfi, qfis, rho0, sld, status = iss_channel_qfi(channel, ancilla_dim)
@@
Fig. 4(B) and 4(C) present plots of the results obtained using iss_channel_qfi and mop_channel_qfi for the two decoherence models respectively. Since the MOP method gives the optimal QFI implicitly assuming arbitrary ancilla dimension $d_{\mathcal{A}}$, these plots indicate that for the dephasing case (B) the ancilla is in fact unnecessary, while in the amplitude damping case (C) the ancillary system of dimension $d_{\mathcal{A}}=2$ is required for an optimal precision.
This example clearly shows the benefits of having two independent optimization procedures. One method (MOP) provides the true optimal QFI, but does not allow us to gain insight into the required ancilla size, while the other one (ISS) allows us to study the impact of the ancilla size.
@@ 3.3 Parallel strategy optimization
Figure 5: (A) Diagram of a parallel strategy with multiple parametrized channels $\Lambda_{\theta}$ and ancilla $\mathcal{A}$. Values of QFI normalized by the number of channel uses, $N$, for: (B) dephasing ($p=0.75$), (C) amplitude damping ($p=0.75$) and different methods: MOP - black $\times$, simple ISS with $d_{\mathcal{A}}=2$ - black +, tensor network ISS with $d_{\mathcal{A}}=2,\;r_{\mathrm{MPS}}=\sqrt{r_{\mathfrak{L}}}=2$ - blue dashed line, tensor network ISS with $d_{\mathcal{A}}=2,\;r_{\mathrm{MPS}}=\sqrt{r_{\mathfrak{L}}}=4$ - red dotted line, upper bound - black solid line.
The parallel strategy refers to a setting in which $N$ copies of $\Lambda_{\theta}$ are probed in parallel by an entangled state and their output is collectively measured, see Fig. 5(A). It can also be understood as a strategy with a single channel $\Lambda_{\theta}^{(N)}=\Lambda_{\theta}^{\otimes N}$. The QMetro++ package provides several functions computing the QFI for the parallel strategy. The simplest two, iss_parallel_qfi and mop_parallel_qfi are straightforward generalizations of ISS and MOP methods for parallel strategy:
@@
8iss_qfi, qfis, rho0, sld, status = iss_parallel_qfi(channel, N, ancilla_dim)
@@
This problem can be circumvented using tensor networks MPS (31) and MPO (32). This approach is implemented in iss_tnet_parallel_qfi, which requires two additional arguments specifying bond dimensions of the MPS and the pre-SLD MPO:
@@
11qfi, qfis, psis, Ls, status = iss_tnet_parallel_qfi(channel, N, ancilla_dim, mps_bond_dim, L_bond_dim)
In contrast to the function implementing the basic ISS, which returns the optimal input state and the SLD in terms of arrays, this procedure gives them in the form of lists containing tensor components: psis = $\psi_{0}$, $\psi_{1}$, …] and Ls = $\mathfrak{L}_{0}$, $\mathfrak{L}_{1}$, …], see Eqs. (31) and (32).
Naturally, increasing ancilla and bond dimensions allows to achieve higher values of QFI up to some optimal value, see Fig. 5(B) and (C). One might assess this value by running iss_tnet_parallel_qfi for progressively larger dimension sizes, but it can also be upper-bounded using the par_bounds function which gives a bound on the QFI of the parallel strategy for all ancilla and bond dimension sizes (see Sec. 2.5):
@@
The numerical results are presented in Fig. 5(B) and (C). We considered the parallel strategy with ancilla dimension $d_{\mathcal{A}}$ and various bond dimensions $r_{\mathrm{MPS}}=\sqrt{r_{\mathfrak{L}}}$. The relation between $r_{\rm{MPS}}$ and $r_{\mathfrak{L}}$ was chosen based on the fact that the density matrix of the input state has bond dimension equal to $r_{\mathrm{MPS}}^{2}$, so taking $r_{\mathfrak{L}}$ up to this value should not significantly affect the execution time (in practice, a smaller $r_{\mathfrak{L}}$ is typically sufficient). Notice that for the dephasing case (B) results from MOP and ISS methods coincide perfectly and that QFI for MPS with $r_{\mathrm{MPS}}=4$ approaches the upper bound as $N$ increases. This suggests that in this case $d_{\mathcal{A}}=2$ is sufficient to saturate the bound. On the other hand, for the amplitude damping case (C) the results obtained using MOP and ISS differ, and already for $N=5$ the ancilla dimension $d_{\mathcal{A}}=2$ is suboptimal. Additionally, $F_{Q}^{(N)}(\Lambda_{\theta})/N$ computed using MPS quickly flattens out and the increase of bond dimension leads to a very small improvement. Therefore, we can conclude that in this case the dimensions $d_{\mathcal{A}}$ considered are insufficient to saturate the bound.
@@ 3.4 Adaptive strategy optimization
Figure 6: (A) Diagram of an adaptive strategy with multiple parametrized channels $\Lambda_{\theta}$ and ancilla $\mathcal{A}$. When the optimization is performed over all quantum combs $C$ (area bounded by dashed lines) one can control the dimension of only the final ancilla ($\mathcal{A}_{\mathrm{end}}$). Values of QFI normalized by the number of channel uses, $N$, for: (B) dephasing ($p=0.75$) (C) amplitude damping ($p=0.75$) and different methods: MOP - black $\times$, simple ISS with $d_{\mathcal{A}}=2$ - black +, tensor network ISS with $d_{\mathcal{A}}=2$ - blue dashed line, tensor network ISS with $d_{\mathcal{A}}=4$ - red dotted line, upper bound - black solid line, asymptotic bound - black dash-dotted line. Lack of data for $d_{\mathcal{A}}=4$, $N\geq 90$ (B) and $N\geq 60$ (C) is due to increasing numerical instability of ISS with growing NN.
In the adaptive strategy, we start with an initial state of the system and the ancilla, $\rho_{0}$, and we act on the system with $N$ parameter-encoding channels $\Lambda_{\theta}$, intertwined with control operations $\mathrm{C}_{i}$ which act both on the system and the ancilla, see Fig. 6(A). Equivalently, it is a strategy where $N$ parameter-encoding channels $\Lambda_{\theta}$ are plugged in between the teeth of a quantum comb $C$, see Fig. 3, and then the resulting state is measured. In this strategy, one seeks to determine the value of QFI for the optimal input state, measurement, and control operations (or equivalently the optimal comb and measurement). Notice that for sufficiently large ancilla dimension $d_{\mathcal{A}}$ and for $\mathrm{C}_{i}$ being appropriate SWAP gates one can simulate any parallel strategy with an adaptive strategy. Thus, assuming the ancilla is large enough, the QFI for the adaptive strategy always upper-bounds the QFI for the parallel strategy.
@@
Alternatively, we may also use the ISS method, with the help of iss_adaptive_qfi function. This procedure will perform the optimization over the whole comb as well. Note that, the constraints on a quantum comb (19) allow for its decomposition into control operations (CPTP maps), but do not specify the ancilla dimension inside the comb. Therefore, in this case one can control only the dimension of the last ancilla which goes outside of the comb, indicated by the symbol $\mathcal{A}_{\mathrm{end}}$ in Fig. 6(A):
@@
10qfi, qfis, comb, sld, status = iss_adaptive_qfi(channel, N, ancilla_dim)
@@
Finally, if we want to make use of the tensor-network structure, and perform the optimization over separate control operations, we can use iss_tnet_adaptive_qfi. Notice that in this approach we control ancilla dimension at each step and thus, in general, the obtained QFI will be smaller than the one we got with iss_adaptive_qfi—the advantage here is that the optimization will be much more efficient (not to mention that this is also the only way to go to the larger $N$ limit):
@@
10qfi, qfis, teeth, sld, status = iss_adaptive_qfi(channel, N, ancilla_dim)
@@
The returned argument teeth represents the comb’s teeth and it is a list consisting of the optimal initial state and CJ matrices of the optimal control operations teeth $\rho_{0},C_{0},C_{1},...,C_{N-2}]$.
The upper bound for the adaptive strategy is computed using ad_bounds function:
@@
Additionally, one can compute the value of the bound in the limit $N\rightarrow\infty$ which is asymptotically saturable and the same for both the parallel and the adaptive strategy, see Sec. 2.5:
@@
6c, k = asym_scaling_qfi(channel)
@@
Values of QFI, bounds, and asymptotic QFI for the adaptive strategy are presented in Figs. 6 (B) and (C). In contrast to the parallel strategy, this time the amplitude damping case (C) shows results which are much closer to the bound.
Notice that there is no data for amplitude damping and ancilla dim $d_{\mathcal{A}}=4$ and $N\geq 60$. This is due to the fact that for this noise model, the QFI attains a large value and simultaneously optimal control operations $\mathrm{C}_{i}$ take the state close to the pure state. These two processes combined create objects (tensors, matrices etc.) which hold values from a very wide range at the same time. This causes loss of precision and then numerical instability. A similar issue appears for dephasing but for much larger $N\simeq 90$.
@@ 3.5 Correlated noise models
Figure 7: (A) Diagram of NN parametrized channels $\Lambda_{\theta}$ with noise correlations. Extended channels $\tilde{\Lambda}_{\theta}$ (gray) act on the system as $\Lambda_{\theta}$ and are connected with each other by the environment space $\mathcal{E}$. The environment system $\mathcal{E}$ starts in the state $\rho_{E}$ and is not accessed (is traced out) at the end. Values of QFI normalized by the number of channel uses, $N$, for anti-correlated (c=0.75c=-0.75) dephasing ($p=0.75$), different strategies: (B) parallel, (C) adaptive and various methods: MOP - black $\times$, simple ISS with $d_{\mathcal{A}}=2$ - black +, tensor network ISS with $d_{\mathcal{A}}=2$ (parallel: $r_{\mathrm{MPS}}=\sqrt{r_{\mathfrak{L}}}=2$) - blue dashed line, upper bound with block size $m=1$ - black dotted line, upper bound with block size $m=3$ - black solid line and upper bound for uncorrelated channels $c=0$ - black dash-dotted line. Notice that since there are no separate bounds for parallel strategy with correlated channels the upper bounds (black dotted and solid lines) are identical in (B) and (C).
The QMetro++ also provides a simple way to compute the QFI for the adaptive and the parallel strategies in scenarios when parameter-encoding channels act on an additional environment space $\mathcal{E}$ which directly connects subsequent channels, see Fig. 7(A). This approach allows in particular to model noise correlations affecting the sensing channels.
To illustrate this, let us consider a Kraus representation for the dephasing model with Kraus operators $K_{\pm}$ acting as Bloch-sphere rotations by $\pm\epsilon$ around $z$ axis, see Eq. (44). We want the signs of these rotations to form a binary Markovian chain. The conditional probability of a rotation with sign $s_{i}\in\{+,-\}$ in channel $i$ given $s_{i-1}$ in channel i1i-1 is:
@@
where $c\in[-1,1]$ is a correlation parameter with $c=0$ meaning no correlations, $c=1$ maximal positive correlations and c=1c=-1 maximal negative correlations. The initial probabilities are $p(\pm)=1/2$.
@@
This process can be modeled with the help of a two-dimensional environment Hilbert space $\mathcal{E}$ which will carry the information about the sign of the rotation in the preceding step. Let us introduce a new channel acting on both the environment and the system $\tilde{\Lambda}_{\theta}=\Phi_{\theta}*\mathrm{T}$, see Fig. 8, where:
[…]
@@
2from qmetro import (
3 par_dephasing, ket_bra, ParamChannel
4)
5
6p = 0.75
7c = -0.75
8
9# rot_like=True gives dephasing in
10# the second form:
11Lambda = par_dephasing(p, rot_like=True)
12krauses, dkrauses = Lambda.dkrauses()
13# Rotations by plus and minus epsilon:
14Kp, Km = krauses
15# Derivatives of Krauses:
16dKp, dKm = dkrauses
17
18# Plus state:
19plus = array([1, 1]) / sqrt(2)
20# Minus state:
21minus = array([1, -1]) / sqrt(2)
22
23# V_theta matrix:
24V = kron(ket_bra(plus, plus), Kp)
25V = V + kron(ket_bra(minus, minus), Km)
26V = sqrt(2) * V
27
28# Derivative of V_theta matrix:
29dV = kron(ket_bra(plus, plus), dKp)
30dV += kron(ket_bra(minus, minus), dKm)
31dV = sqrt(2) * dV
32
33# env_dim=d tells ParamChannel
34# constructor that the channel acts
35# on the d-dimensional environment space
36# and the system
37Phi = ParamChannel(